# Liwen Wang1,2,  Yuanyuan Yuan2,  Ao Sun2,  Zongjie Li2,  Pingchuan Ma2,  Daoyuan Wu2,  Shuai Wang2

### v2

@@ 3.3 MLLM-Based Mutations
Instruction-Driven. MLLM’s success in the multimodal domain has revolutionized the image mutation paradigm. MLLMs enable translation between text and images, such that users can freely describe their anticipated mutation to a seed image using text instruction; MLLMs then generate the mutated image according to the instruction. This instruction-driven paradigm largely expands the covered semantics in image mutations and liberates the users’ freedom to design the anticipated mutations. Moreover, by designing the proper text instruction, previous mutations discussed in Sec. 3.1 and Sec. 3.2 can be uniformly implemented using MLLMs.
Semantic-Additive. Besides the unified implementation of mutations mentioned above, the instruction-driven paradigm also introduces a new category of mutations, which we refer to as semantic-additive mutations. Instead of modifying existing semantics in images, these mutations can bring new semantics to the seed image (which are nearly infeasible using previous mutations). For example, the dog image in Fig. 1 can be decorated with clothes to challenge the VDL system under test. Essentially, such additive capability enables a new dimension for the input mutation in VDL testing.
Knowledgeable readers may wonder the differences between semantic-additive mutations with the object insertion adopted in testing object-based VDL systems [57, 66, 64]. We clarify that semantic-additive mutations operate in semantic-level and are finer-grained, which often require reasoning the relation and hierarchy between semantics (either intra- or inter-object). For example, the “dressing with clothes” mutation in Fig. 1 needs to understand and identify the dog’s body and replace it with appropriate clothes. Object insertion, in contrast, does not perceive image semantics and only relies on object locations in the image, so that the inserted object will not overlap with existing ones. Since object insertion mostly applies to object-based VDL systems (e.g., the inserted object may affect the classification result), we omit it in this study.
@@ 4 Research Motivations
Limitations of Past Methods. Though the fast development of input mutations in VDL testing has achieved fine-grained control over higher-level image semantics, they only apply to limited scenarios and specific seeds. Noteworthy, existing perceptual-level mutations (the most recent and effective mutation) require humans to annotate the mutated semantics in each seed and let humans validate and interpret the achieved mutation [52, 71]; this process only applies to highly aligned images (e.g., face photos taken in a front view where a unified mask can be adopted to pinpoint eyes, mouth, etc.) and can hardly be automated in practice. From this perspective, it is expected that MLLMs can offer a more general, automated, and highly flexible mutation scheme for VDL testing. The mutations can be controlled by text instructions, which are more intuitive, accessible, and calibratable to users than past methods.
Unclear MLLM Internals. Despite the optimism, the internals of how such mutations are achieved by MLLMs remain unclear. Previous explicit mutations (pixel- and geometrical-level mutations mentioned in Sec. 3.1) are rigorously implemented according to their mathematical definitions. Implicit mutations (style- and perceptual-level mutations in Sec. 3.2) often rely on well-developed generative models to ensure generating valid images; the usefulness of these methods has been empirically justified in recent works [35, 16, 65].
In contrast, per recent reports, MLLMs often fail to mutate images according to the given instructions [41], and may generate broken and ill-formed images from time to time [18, 10, 21]. Hence, we believe our community still lacks understanding of how exactly MLLMs can boost VDL testing, and a comprehensive study is urgent for VDL developers to understand MLLM’s (in-)capability of mutating images, especially in the context of VDL testing where the quality of mutated test inputs is crucial.
@@ 5 Research Study Setup
Our key research question (RQ) is on MLLM’s capability of generating test inputs for VDL testing. To this end, we conduct both human studies and quantitative evaluations to assess if MLLM can support varying mutations noted in Sec. 3. The setups are introduced below.
@@ 5.1 Evaluated Aspects
Following Sec. 4, we focus on four aspects to assess the quality of image mutations (both traditional and MLLM-enabled) for VDL testing. These four aspects include three input-side requirements, and one output-side assessment.
@@
Faithfulness. A mutation should faithfully preserve the semantics that are ought to be unchanged. For example, to close eyes of face photos, the mutation should let other semantics (e.g., background, hair, nose) remain unchanged and only close the eyes. Otherwise, the unanticipated mutation may break the testing oracle and brings false positives. The follow-up diagnosis for VDL’s faulty behaviors can also be misled.
Validity. Different from traditional software (e.g., a sorting algorithm) that implement input checker like type checker, the deep learning module in modern VDL system often does not support checking whether an input image is semantically valid. Since input images are represented as numerical matrices, the VDL system can evoke all computations and generate a valid, but random output when taking an image containing only noise or a fragmented image.
Overall, from a high level, alignment focuses on the mutated semantics whereas faithfulness asserts the unchanged semantics. Validity more holistically assesses the semantics in the whole image.
Testing Effectiveness. Further to the above three input-side requirements, we also consider the testing effectiveness as an output-side assessment: we check how well the mutated images can detect VDL faults. This is crucial as we need to assess whether a mutation is valuable. The MLLM’s instruction-driven paradigm is free-form, meaning that innumerous mutations can be (potentially) enabled. For effective and efficient testing, users require assessing which mutations are more useful in detecting VDL faults.
@@ 5.2 MLLMs and Their Pipelines
We evaluate four SoTA and representative MLLMs that support both text and image inputs, including LLaVA [43], CogVLM [58], GPT-4V [5], and InstP2P [8].
@@
Dialog-Based. The most straightforward way to mutate images is establishing a dialog with the MLLM, as shown in Fig. 2(a). For example, the users first provide the seed image and a prompt of the mutation instruction to GPT-4V and let GPT-4V return the mutated image. However, our preliminary explorations show that this dialog-based pipeline fails to generate test inputs in all cases we try: GPT-4V either generates an entirely new image (nearly all semantics are changed) or fails to achieve the mutation (i.e., GPT-4V responds that it cannot generate the image).
@@
To alleviate this issue, one may expect using prompt engineering to better guide the MLLM. Nevertheless, our preliminary study shows that it is hard to tune MLLM prompts and achieve constantly improved performance. Moreover, we argue that “tuning prompts” is less practical in our context, as VDL developers (main users of VDL testing tools) are often not experts in prompt engineering; MLLM should provide out-of-the-box support for implementing mutations. Therefore, following suggestions in [22, 33, 11, 9, 32], we consider optimizations over the mutation pipeline and equip the MLLM with Diffusion model [4], a more sophisticated image generation model, as introduced below.
Figure 2: Pipelines of MLLM-based mutations.
@@
Unidirectional & Fine-tuning-based. InstP2P designs an unidirectional pipeline (Fig. 2(b)) which directly uses a text instruction (the MLLM is employed to extract text features) and the seed image to generate the mutated image. In general, since the text instruction primarily describes the mutation and does not include details in the images, InstP2P fine-tunes the pre-trained Diffusion model for a diverse set of mutations. Each fine-tuning data point consists of a seed image and its mutated variant, and the Diffusion model is trained to generate the mutated image when given the seed image and the corresponding text instruction. We refer interested readers to [8] for the fine-tuning details.
@@
Bidirectional & Post-hoc. Another pipeline, as designed in [22, 33, 11, 9, 32], leverages MLLMs to transform the mutation instruction based on the seed image, such that details in the seed image can be included in the transformed instruction. This pipeline is categorized as bidirectional (Fig. 2(c)) as it implements the workflow of image $\rightarrow$ text $\rightarrow$ image. Specifically, due to MLLM’s promising capability of understanding images, users first feed the seed image to the MLLM and let the MLLM describe what the image would look after applying a mutation. Here, the MLLM’s output is a textual description of semantic-level details in the mutated image. This finer-grained text description, together with the seed image, are then fed to the Diffusion model to generate the mutated image. The superiority of this pipeline (compared with the dialog-based one) has been empirically justified in [22, 33, 11, 32].
@@
Overall, the bidirectional pipeline, to some extent, achieves an automated prompt engineering by enriching the mutation instruction with the seed image’s details. The unidirectional pipeline, on the other hand, fine-tuning the Diffusion model to strengthen its capability in achieving image mutations. In our evaluations, we implement CogVLM, LLaVA, and GPT-4V with the bidirectional pipeline, and use the original unidirectional pipeline for InstP2P.
@@ 5.3 Datasets and Mutations
Table 1 lists the evaluated datasets. They are large-scale and representative to cover different application scenarios. ImageNet consists of 1,000 classes of real-life images, and is widely used for general image classification tasks. Dog-Breed dataset contains dog images with 120 different breeds and is a subset of ImageNet. It is constructed to improve VDL model’s accuracy in domain-specific applications. FFHQ includes human face photos; it is widely used for face recognition and annotation tasks. CityScape is composed of urban street scene images, which are popular in autonomous driving, object detection, and instance segmentation, etc.
Table 1: Evaluated datasets.
@@
ImageNet [14] 1,000 Diverse real-life images Image Classification
Dog-Breed [2] 120 Dog images of diff. breeds Image Classification
FFHQ [36] N/A Human face photos Face recognition, annotation, etc.
@@
Table 2: Evaluated mutations and their text instructions when implemented with MLLMs.
@@
② Rotation Geometrical Rotate the image 180 degrees.
@@
④ SnowG What would it look like if it were snowing.
⑤ SnowD What would it look like if it were snowing.
⑥ Eyes Perceptual Make the eyes of the person close.
⑦ Background Semantic- additive Change the background to a library.
@@
Aligned to different types of mutations introduced in Sec. 3, we consider six mutations (covering all four types) supported by traditional methods, as shown in Table 2. Specifically, since the weather transfer mutation is affected by the to-be-mutated image, we evaluate it on both general images (SnowG in Table 2) and driving scene photos (SnowD in Table 2) from ImageNet and CityScape, respectively. The perceptual-level mutation is achieved using GANs [71] and is dataset-specific, which only applies to aligned face photos in FFHQ. We also consider four different semantic-additive mutations that are uniquely enabled by MLLM. They operate in different granularities and manifest varying characteristics, as implied in the last column of Table 2. Considering that ⑧-⑩ require specifying the mutated object in the image, to evaluate them uniformly on diverse images, we use the Dog-Breed dataset.
To implement a mutation (either traditional or MLLM-enabled) using MLLMs, a text instruction (i.e., prompt) is required. We list the text instruction of each mutation in Table 2. Recall as mentioned in Sec. 5.2, we consider two optimized MLLM pipelines that are specifically designed for achieving image mutations, so that a mutation’s prompt can be a single-sentence instruction, enabling out-of-the-box usage. Due to such a simple form, we find that the MLLM is mostly robust to changes over those instructions, and we observe that, the instruction’s wording, grammatical structure, etc., does not notably impact the achieved mutation.
Table 3: Results of human evaluations. We mark results within $[0,1)$, $(2,4]$, and $(4,5]$ in red, blue, and green, respectively.
@@
  • Green: Perfectly satisfy. Blue: Satisfy. Red: Fail.
6 Human Studies

### v3

@@ 3.3 MLLM-Based Mutations
Unification & Generalization: Instruction-Driven. MLLMs have revolutionized image mutation by enabling text-to-image translation. Users can describe the desired mutation for a seed image using text instructions, and MLLMs generate the mutated image accordingly. This instruction-driven paradigm enables VDL developers to uniformly implement previous mutations introduced in Sec. 3.1 and Sec. 3.2 and apply domain-specific mutations in Sec. 3.2 to more general scenarios, by designing the proper text instruction.
Expansion: Semantic-Replacement. The instruction-driven paradigm also largely expands the covered semantics in image mutations. Unlike previous mutations that edit the status of semantics already presented in the seed image, MLLMs bring new mutations that can replace image semantics with new ones; we refer to this new category as semantic-replacement in the rest of this paper. For example, the dog image’s main body can be replaced with clothes (see Fig. 1) to further challenge the VDL system’s recognition.
Comparison with Object Insertion. Knowledgeable readers may wonder the differences between semantic-replacement mutations with the object insertion adopted in testing object-based VDL systems [2, 36, 3]. We clarify that semantic-replacement mutations operate in semantic-level and are finer-grained, which often require reasoning the relation and hierarchy between semantics (either intra- or inter-object). For example, the “dressing with clothes” mutation in Fig. 1 needs to understand and identify the dog’s body and replace it with appropriate clothes. Object insertion, in contrast, does not perceive image semantics and only relies on object locations in the image, so that the inserted object will not overlap with existing ones. Moreover, prior object insertion mutation is not end-to-end, as they rely on object localization results from object detectors; we thus omit it in this study.
@@ 4 Research Motivations
Limitations of Past Methods. Although the fast development of input mutations shows promising results in VDL testing, maintaining and developing them often requires expertise in different domains. Moreover, the more advanced style- and perceptual-level mutations show limited application scopes. For instance, weather-targeted mutations may perform worse in non-driving-scene images, and the face-related perceptual-level mutations only apply to highly aligned face photos (as they rely on a unified mask to pinpoint eyes, mouth, etc.). Some perceptual-level mutations may require manually annotating the to-be-mutated semantics [9, 13].
From these perspectives, we envision that MLLMs can offer an instruction-driven mutation paradigm, which is more general, automated, and highly flexible for VDL testing. MLLMs have shown high capabilities in natural language understanding and image synthesis, and therefore, various kinds of mutations that were hardly or tediously achieved by traditional methods may be easily implemented via MLLMs by directly providing instructions. Moreover, since various (subtle or complex) mutations are unifiedly described in natural language (as shown in Fig. 1), the mutations can be easily understood, maintained, calibrated, and extended by VDL developers and the community. [^4] In some sense, this new instruction-driven mutation paradigm is comparable to how query-based software security testing tools, e.g., CodeQL [52], maintain varying types of vulnerabilities in a unified query language.
Unclear MLLM Internals. Despite the optimism, MLLM’s image generation is hard to interpret due to the unclear internals. Per recent reports, MLLMs often fail to mutate images according to the given instructions [53], and may generate broken and ill-formed images from time to time [27, 28, 29]. Hence, we believe our community still lacks understanding of how exactly MLLMs can boost VDL testing, and a comprehensive study is urgent for VDL developers to understand MLLM’s (in-)capability of mutating images. In addition, considering that input validation is crucial to ensure meaningful testing results, it is also demanding to evaluate whether existing validation metrics can still be applied to MLLM-based mutations.
@@ 5 Research Study Setup
Our study answers the following research questions (RQs).
  • RQ1: Can MLLMs unify different traditional input mutations? We answer this RQ by assessing the quality of 1) how MLLMs re-implement traditional mutations and 2) how MLLMs extend prior domain-specific mutations to general scenarios.
  • RQ2: What benefits can MLLMs bring to input mutations in VDL testing? We answer this RQ by 1) assessing the quality of test inputs generated by the semantic-replacement mutations, and 2) evaluating their effectiveness in identifying VDL faults.
  • RQ3: Are existing input validation metrics reliable when facing MLLM-based mutations? We answer this RQ by evaluating whether existing validation metrics can reflect 1) the validity of test inputs generated by MLLMs and 2) the varied extents of different MLLM-based mutations.
Our RQs are based on the quality assessment of test inputs generated via different mutations, which requires perceiving and understanding image semantics. We therefore conduct a large-scale human study to evaluate image quality. Below, we introduce our evaluated aspects for the quality assessment.
@@ 5.1 Evaluated Aspects
Following design principles in existing input mutations and validation metrics, we consider the following three aspects.
@@
Faithfulness. A mutation should faithfully preserve the semantics that are ought to be unchanged. For example, to close eyes of face photos, the mutation should let other semantics (e.g., background, hair, nose) remain unchanged and only close the eyes. Otherwise, the unanticipated mutation may break the testing oracle and bring false positives. The follow-up diagnosis for VDL’s faulty behaviors can also be misled.
Validity. Besides the alignment and faithfulness that focus on the mutated and unmutated semantics, we also holistically consider the validity of all semantics in the whole image. This is crucial as an aligned and faithful mutation may still generate unrealistic images. For example, as will be shown in Fig. 4, GPT-4V replaces a portrait’s eyes with closed ones (potentially from other portraits in its “database”) when asked to close the eyes. Although the mutation aligns the requirement and other semantics are untouched, the sharp transition around eyes makes the mutated image unrealistic, which rarely appear in real-world scenarios.
@@ 5.2 MLLMs and Their Pipelines
MLLMs. To date, two pipelines are proposed by existing works to employ MLLMs for image mutations. As illustrated in Fig. 2(b)-(c), these pipelines are optimized from different angles to avoid issues in prompt engineering or image generation modules, providing out-of-the-box usages for non-expert in MLLMs. In particular, the first pipeline (Fig. 2(b); see more details below) relies on a specifically fine-tuned MLLM, we therefore use the MLLM provided by the authors [54], along with this pipeline, as a whole for evaluations and refer to it as InstP2P. The second pipeline is post-hoc and supports incorporating different MLLMs; thus, we use three SoTA and representative MLLMs, including LLaVA [55], CogVLM [56], and GPT-4V [17] into this pipeline during evaluations.
Figure 2: Pipelines of MLLM-based mutations. Fig. 2(a) shows the most straightforward dialog-based pipeline. Fig. 2(b)-(c) illustrates two pipelines specifically optimized for image mutations. The unidirectional pipeline shown in Fig. 2(b) requires fine-tuning a MLLM and should be used along with this tailored MLLM. The bidirectional pipeline presented in Fig. 2(c) is post-hoc and supports incorporating different MLLMs.
@@
Dialog-Based. The most straightforward way to mutate images is building a dialog with the MLLM, as shown in Fig. 2(a). For example, the users first provide the seed image and a prompt of the mutation instruction to GPT-4V and let GPT-4V return the mutated image. However, our preliminary explorations show that this dialog-based pipeline fails to generate test inputs in all cases we try: GPT-4V either generates an entirely new image (nearly all semantics are changed) or responds that it cannot generate the image. Similar issues are also noted in [57, 58, 59],
@@
To alleviate this issue, one may expect to use prompt engineering to better guide the MLLM. Nevertheless, our preliminary study shows that it is hard to automatically tune MLLM prompts and achieve improved performance constantly on different images. Moreover, we argue that “tuning prompts” is less practical in our context, as VDL developers (main users of VDL testing tools) are often not experts in prompt engineering; MLLM should provide out-of-the-box support for implementing mutations. In this regard, recent works propose to equip MLLMs with Diffusion model [60], a more sophisticated image generation model, and design two representative optimization pipelines to orchestrate their specialties in cross-modal translation and image generation [57, 58, 61, 62, 59].
@@
Unidirectional & Fine-tuning-based. As shown in the bottom of Fig. 2(b), InstP2P designs an unidirectional pipeline which directly uses a free-form text instruction and the seed image to generate the mutated image. In general, since the text instruction primarily describes the mutation and does not include details in the images, InstP2P requires fine-tuning a Diffusion model to support diverse mutations.
Specifically, as illustrated in the top and middle of Fig. 2(b), a GPT model is first employed to edit a given textual caption to obtain a new caption. Then, the two captions are fed into a pretrained a Stable Diffusion model [60] to generate two images, $i_{s}$ and $i_{t}$, respectively. Finally, a free-form textual instruction is generated using GPT to briefly summarize the changes between the two captions; this textual instruction, together with the two generated images $i_{s}$ and $i_{t}$, forms a data point for fine-tuning the Diffusion model, and the Diffusion model is tuned to generate $i_{s}$ when given $i_{t}$ and the textual instruction. After fine-tuning, the Diffusion model is able to generate a mutated image for any unseen images and textual instructions. We refer interested readers to [54] for the fine-tuning details.
@@
Bidirectional & Post-hoc. Another pipeline, as designed in [57, 58, 61, 62, 59], leverages MLLMs to transform the mutation instruction based on the seed image, such that details in the seed image can be included in the transformed instruction. As shown in Fig. 2(c), this pipeline is categorized as bidirectional, since it implements the workflow of image $\rightarrow$ text $\rightarrow$ image. Specifically, due to MLLM’s promising capability of understanding images, users first feed the seed image to the MLLM and let the MLLM describe what the image would look like after applying a mutation. Here, the MLLM’s output is a textual description of semantic-level details in the mutated image. This finer-grained text description, together with the seed image, is then fed to the Diffusion model to generate the mutated image. The superiority of this pipeline (compared with the dialog-based one) has been empirically justified in [57, 58, 61, 59].
Fig. 3 shows examples of textual descriptions in the bidirectional pipeline. The text can precisely describe the mutation expected by the instruction. For instance, in the first case shown in Fig. 3, the description “The lizard would be surrounded by a snowy environment…” precisely captures the semantic transformation specified by the mutation instruction “What would it look like if it were snowing”, facilitating accurate execution of the intended mutation. In addition, the text includes many detailed and informative descriptions, e.g., “The snow would also cover the tree branch and the surrounding area…” in the first case, which gives more information of the mutated image to ease the mutant generation.
Figure 3: Textual descriptions generated in the bidirectional & post-hoc pipeline of MLLM-based mutations.
@@
Overall, the bidirectional pipeline, to some extent, achieves automated prompt engineering by enriching the mutation instruction with the seed image’s details. The unidirectional pipeline, on the other hand, fine-tuning the Diffusion model to strengthen its capability in achieving image mutations. In our evaluations, we incorporate three MLLMs, namely, CogVLM, LLaVA, and GPT-4V, into the bidirectional pipeline, resulting in three MLLM-based mutation methods. We use the unidirectional pipeline with its tailored MLLM and denote this mutation method as InstP2P. It’s worth noting that CogVLM, LLaVA, GPT-4V, and InstP2P (before fine-tuning) use the same Stable Diffusion model [60] as the image generation module, ensuring a fair comparison for the capabilities of different MLLMs.
@@ 5.3 Datasets and Mutations
Table I lists the evaluated datasets. They are large-scale and representative to cover different application scenarios. ImageNet consists of 1,000 classes of real-life images, and is widely used for general image classification tasks. Dog-Breed dataset contains dog images with 120 different breeds and is a subset of ImageNet. It is constructed to improve VDL model’s accuracy in domain-specific applications. FFHQ includes human face photos; it is widely used for face recognition and annotation tasks. CityScape is composed of urban street scene images, which are popular in autonomous driving, object detection, and instance segmentation, etc.
TABLE I: Evaluated datasets.
@@
ImageNet [31] 1,000 Real-life images Image Classification
Dog-Breed [32] 120 Dog images Image Classification
FFHQ [33] N/A Face photos Face recognition, annotation, etc.
@@
TABLE II: Mutations and their text instructions in MLLMs.
@@
② Rotation Geo. Rotate the image 180 degrees.
@@
④ SnowG What would it look like if it were snowing?”(for general images)
⑤ SnowD What would it look like if it were snowing?”(for driving scene photos)
⑥ Eyes Percep. Make the eyes of the person close.
⑦ Background MLLM- enabled Change the background to a library.
@@
  • SnowG: weather transfer mutations applied on general images. SnowD: weather transfer mutations applied on driving scene photos.
Traditional. We consider five representative mutations for the five categories of traditional mutations introduced in Sec. 3. Additionally, since the weather transfer mutation is affected by the to-be-mutated image, we evaluate it on both general images (SnowG in Table II) and driving scene photos (SnowD in Table II) from ImageNet and CityScape, respectively. The perceptual-level mutation is achieved using GANs [13]; it is dataset-specific and only applies to aligned face photos.
MLLM-Enabled. For the semantic-replacement mutations uniquely enabled by MLLMs, we consider four different variants operating in different granularities and manifesting varying characteristics. Background only edits background semantics, while the remaining three focus on finer-grained semantics in the main subject. Glasses is more “silent”: it replaces the original eyes with glasses + new eyes, and the new eyes should be mostly similar to the original ones, with only slight updates on the lighting condition for coherence. Clothes and Legs, in contrast, are more active by completely replacing target semantics with new ones, and their target semantics are of interest to different VDL models when recognizing the dog. Since these four mutations require specifying the mutated object in the images, to evaluate them uniformly on diverse images, we use the Dog-Breed dataset. That said, extending them to other objects/images should be straightforward: users only need to replace the word “dog” with the target one.
Prompt. To implement a mutation (either traditional or MLLM-enabled) using MLLMs, a text instruction (i.e., prompt) is required. We list the text instruction of each mutation in Table II. Recall as mentioned in Sec. 5.2, we consider two optimized MLLM pipelines that are specifically designed for achieving image mutations, so that a mutation’s prompt can be a single-sentence instruction, enabling an out-of-the-box usage. Due to such a simple form, we find that the MLLM is mostly robust to changes over those instructions, and we observe that, the instruction’s wording, grammatical structure, etc., does not notably impact the achieved mutation. We thus use the same instructions for both MLLM pipelines in our evaluations; this setup also allows us to evaluate the specialties of different pipelines and motivates future pipeline design.
TABLE III: Results of human evaluations. We mark results within $[0,1)$, $(2,4]$, and $(4,5]$ in red, blue, and green, respectively.
@@
6 Human Studies Setup